# TODO: Many title candidates are possible, but only the last taken

### v1

@@ 1 Introduction
Remote sensing applications often grapple with the challenge of processing vast amounts of data distributed across various geographical locations. As the demand for sophisticated image classification models grows, the need for efficient and privacy-preserving methodologies becomes imperative. If we collect data in remote sensing and train a centralized model, we will encounter privacy issues. On the contrary, if we train a local model, we will face challenges related to non-uniform distribution and insufficient data across devices.
Federated Learning (FL) emerges as a promising solution, allowing decentralized devices to collaboratively train machine learning models while keeping raw data localized. FL has various approaches such as FedAVG, FedSGD, FedBN, etc. [1] and can even adopt a peer-to-peer (P2P) [2] model instead of the traditional client-server architecture. FedAVG [3], or Federated Averaging, has been a key algorithm in the domain of FL. It enables multiple decentralized devices or clients to collaboratively train a global model without sharing raw data. This approach ensures privacy by keeping raw data localized while allowing collaborative model training across a distributed network. FedAVG is renowned for its efficiency, simplicity and longevity, acting as a baseline for other algorithms to develop.
Remote sensing problems usually rely on devices that may be located far from the ground, limiting the information being communicated [4]. Thus, there is a crucial need for FL methods that function on small amount of information. In this context, FedAVG has been primarily adopted [5] for aggregating models across devices.
In this paper, we explore various FL communication strategies for remote sensing image classification based on the well-known FedAVG algorithm and study the trade-off between reducing data communicated and performance. In particular, we propose sending and receiving only the average feature vectors for each categories of interest which incurs minimal information. As such, one device could gain insights from the features extracted by the other clients without explicitly synchronizing the parameters between separated training device. The idea that a model can learn from feature representations instead of using network parameters has been widely applied between images and text [6] as CLIP [6] employs Vision Transformer (ViT) to learn feature representations of text extracted from BERT, aiming to maximize the similarity of (image, text) pairs using a contrastive loss. To that end, the Large Margin Cosine Loss [7] is employed to solve the data imbalance problem and allow faster training convergence. Regarding to improve network performance, we combine the FedAVG algorithm with average feature communication and formulate classification as a retrieval problem, obtaining the best of both worlds, high performance with fast convergence.
@@ 2.2 Feature-based FedAVG
This method, however, may result in each client learning diverging representations from one another due to non-uniform variation of the local datasets and random weight initialization. To constrain the deviation of the extracted features a regularization method is introduced. In particular, the same set of random initial parameters, denoted by $M_{0}$ is used to initialize all clients’ backbones on the first round and modulate the backbones’ learned parameters of each client as follows:
@@
where $M_{i}$ is the backbone’s set of parameters of a client after round $i$ and $\tilde{M}_{i}$ is the parameter set updated from the backpropagation. Equation 5 is used after Algorithm 2 before starting a new round. Subsequently, all the clients start from the same set of random parameters $M_{0}$ and the newly learned parameters are scaled down by $\frac{1}{n_{\text{client}}}$. This will be referred to as Algorithm 3.
@@ 2.4 Retrieval-based classification
The idea requires a preparation step before prediction can be made, which includes re-passing all training images through each client’s backbone to extract the feature vectors, similar to step 3 in Algorithm 2, which are distributed to all other clients via the server. Each client classifies an input image by extracting its feature using the backbone model and retrieving the label of the majority of the nearest training features in the embedding space.
@@ 3.1 Datasets and Setup
In this work, we use two remote sensing scene classification datasets to conduct our experiments, namely the UC-Merced (UCM) [11] containing 100 images of size $256\times 256$ per class with 21 land-use classes and the Aerial Image Dataset (AID) [12] with $10,000$ images of size $600\times 600$ over $30$ scene classes. We use the UCM dataset for quick evaluation and AID for more analyses. The Scikit-learn111[Online] Scikit-learn. URL https://scikit-learn.org/. Visisted on 10 Jan, 2024.’s train_test_split function is used to split the dataset into training and validation set with test_size of 0.3.
@@ 3.2 Results
Figure 2 shows the results obtained for the UCM dataset with different FL strategies. The non-Fed approach where the clients are trained independently without any communication is shown for reference purpose. The two feature-based Algorithms 2 and 2.2 converge more rapidly than the baseline FedAVG (blue) yet under-perform it in the long run. The superior of Algorithm 2.2 over 2 shows the benefit of regularization. Combining feature and model-based communication benefits Algorithm 4 and 5 in both ways, surpassing the baseline and maintaining till the final round. The retrieval method for obtaining predicted labels in Algorithm 5 show further performance improvement.

### v2

@@ 1 Introduction
Remote sensing applications often deal with the challenge of processing vast amounts of data collected from, and thus distributed across geographical locations. Locally rounding up collected data for training a centralized model may infringe privacy regulations while training separate local models results in sub-optimal performance due to geographically non-uniform data distribution and overfitting. As the demand for sophisticated image classification models grows, the need for efficient and privacy-preserving methodologies becomes imperative.
Federated Learning (FL) emerges as a potential solution to enforce privacy by allowing decentralized devices to collaboratively train machine learning models across a distributed network while keeping raw data localized. Several approaches such as FedAVG, FedSGD, FedBN, etc. [1] and even a peer-to-peer (P2P) [2] model have been proposed for FL, among which, FedAVG [3], or Federated Averaging, is well-known for its efficiency and simplicity, acting as a baseline for other algorithms. FedAVG has also been adopted for aggregating models across devices for remote sensing [4] as remote sensing tasks are usually run on devices located far from the ground, limiting the information being communicated [5], demanding that FL methods operating on small amount of information.
In this paper, we explore various FL communication strategies for remote sensing image classification based on the well-known FedAVG algorithm and study the trade-off between reducing data communicated and performances. In particular, we propose sending and receiving only the average feature vectors for each category of interest, thus incurring minimal exchanged information. As such, one client could gain insights from the features extracted by the other clients without explicitly synchronizing the trained weights among them. The idea that a model can learn from feature representations instead of using network parameters has been widely applied between images and text as CLIP [6] employs Vision Transformer (ViT) to learn feature representations of text extracted from BERT, aiming to maximize the similarity of (image, text) pairs using a contrastive loss. To that end, the Large Margin Cosine Loss [7] is employed to solve the data imbalance problem and allow faster training convergence. In regard to improving network performance, we combine the FedAVG algorithm with average feature communication and formulate the classification problem as a retrieval problem, obtaining the best of both worlds, high performance with fast convergence.
@@ 2.2 Feature-based FedAVG
This method, however, may result in each client learning diverging representations from one another due to non-uniform variations of the local datasets and random weight initialization. To constrain the deviation of the extracted features a regularization method is introduced. In particular, the same set of random initial parameters denoted by $M_{0}$ is used to initialize all clients’ backbones on the first round and modulate the backbones’ learned parameters of each client as follows:
@@
where $M_{i}$ is the backbone’s set of parameters of a client after round $i$ and $\tilde{M}_{i}$ is the parameter set updated from the backpropagation. Equation 5 is used after Algorithm 2, before starting a new round. Subsequently, all the clients start from the same set of random parameters $M_{0}$ and the newly learned parameters are scaled down by $\frac{1}{n_{\text{client}}}$, which will be referred to as Algorithm 3.
@@ 2.4 Retrieval-based classification
The idea requires a preparation step before prediction can be made, which includes re-passing all training images through each client’s backbone to extract the feature vectors, similar to step 3 in Algorithm 2. The vectors are distributed to all other clients via the server. Each client classifies an input image by extracting its features using the backbone model and retrieving the majority label of the nearest training features in the embedding space.
@@ 3.1 Datasets and Setup
In this work, we use two remote sensing scene classification datasets to conduct our experiments, namely the UC-Merced (UCM) [11] containing 100 images of size $256\times 256$ per class with 21 land-use classes and the Aerial Image Dataset (AID) [12] with $10,000$ images of size $600\times 600$ over $30$ scene classes. We use the UCM dataset for quick evaluation and AID for more analyses. The Scikit-learn111[Online] Scikit-learn. URL https://scikit-learn.org/. Visisted on 25 May, 2024.’s train_test_split function is used to split the dataset into training and validation set with test_size of 0.3.
@@ 3.2 Results
Figure 2 shows the results obtained for the UCM dataset with different FL strategies. The non-Fed approach where the clients are trained independently without any communication is shown for reference purposes. The two feature-based Algorithms 2 and 2.2 converge more rapidly than the baseline FedAVG (blue) yet under-perform it in the long run. The superior of Algorithm 2.2 over 2 shows the benefit of regularization. Combining feature and model-based communication benefits Algorithm 4 and 5 in both ways, surpassing the baseline and maintaining till the final round. The retrieval method for obtaining predicted labels in Algorithm 5 show further performance improvement.